# Fine-Grained Perturbation Guidance via Attention Head Selection

### v3

@@ 1 Introduction
Diffusion models ho2020ddpm ; Song2020ddim ; jascha2015 ; yang2019estgrad ; yang2020sde ; dhariwal2021diffusion ; rombach2022high ; dustin2023sdxl and flow-matching models Lipman2022FlowMatching ; nanye2024sit ; esser2024scaling have gained great popularity in visual generation tasks, including images rombach2022high ; dustin2023sdxl ; peebles2023scalable ; chen2024pixart ; esser2024scaling , videos ho2022video ; Ho2022ImagenVH ; Blattmann2023StableVD ; Zhuoyi2025cogvideox , 3D poole2022dreamfusion ; lin2023magic3d ; wang2023prolificdreamer ; shi2023mvdream , and 4D content Singer2023TextTo4D ; Yu20244Real ; Rundi2025cat4d . The key behind the success of diffusion models lies in classifier-free guidance (CFG) ho2022classifier , which substantially enhances image generation quality during conditional inference. Despite its effectiveness, CFG has two key limitations. First, it applies only to conditional generation, limiting its applicability in unconditional settings such as inverse problems zahra2021linear ; Chung2022DiffusionPS ; Song2023Pseudoinverse ; Chung2022ImprovingDM . Second, CFG often reduces sample diversity and leads to over-saturated or overly simplified outputs seyedmorteza2023cads ; Kynkäänniemi2024Applying ; Saharia2022PhotorealisticTD ; tero2024 .
To address the limitations of CFG in unconditional scenarios, alternative guidance strategies ahn2024self ; hong2023improving ; hong2024smoothed ; tero2024 ; hyung2025spatiotemporal have been proposed. These methods steer the denoising trajectory by perturbing the input or the model itself, or by training a weaker model, thereby guiding samples away from low-quality regions and toward the high-quality data manifold. Among these strategies, attention-layer perturbation approaches ahn2024self ; hong2024smoothed continue to be a practical and widely explored strategy, as they can generate well-aligned weak models without additional training. Some papers explain these guidance methods from an energy perspective hong2024smoothed and draw connections to Hopfield networks demircigil2017model ; ramsauer2020hopfield , which empirically work well with specific blocks (medium blocks of U-Net ronneberger2015u architecture) of attention.
However, there is still limited understanding of where perturbations should be applied. In particular, Diffusion Transformer (DiT) peebles2023scalable architectures lack localized blocks responsible for global semantics, unlike U-Nets ronneberger2015u , and instead distribute this functionality more evenly across all layers avrahami2024stable . This structural difference makes it even more critical to carefully select perturbation targets to achieve effective guidance.
To determine suitable perturbation targets, we first consider the underlying structure of DiT. In DiT, multi-head self-attention vaswani2017attention plays a central role in modeling global dependencies. Each head attends to different aspects of the input, and prior work has shown that heads often specialize in distinct semantic features vaswani2017attention ; park2024cross ; dosovitskiy2020image ; gandelsman2023interpreting . This suggests that attention heads, rather than entire layers, may serve as more precise and effective targets for perturbation.
@@
Building on these observations, we (i) analyze key properties of head-level perturbation guidance, including composability and controllability, (ii) propose HeadHunter, a systematic framework for retrieving heads that align with arbitrary objectives, and (iii) introduce SoftPAG, a variant of PAG ahn2024self that enables fine-grained, continuous control over guidance strength, mitigating over-smoothness and oversimplification caused by overly aggressive perturbations.
@@ 2 Preliminaries
2.1 Generative Modeling Frameworks
Diffusion models.
Diffusion models ho2020ddpm ; sohl2015deep ; song2020score define a generative process as iterative denoising of a sample drawn from a Gaussian prior. The forward process gradually perturbs a clean data point $\mathbf{x}_{0}\sim p_{\text{data}}$ through the noise schedule $(\alpha_{t},\sigma_{t})$:
$\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\quad t\in[0,1].$ (1)
A neural network $\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)$ learns to predict the added noise $\boldsymbol{\epsilon}$, enabling reverse-time sampling from noise to data.
Flow matching.
Flow matching lipman2022flow ; liu2022flow provides a deterministic alternative by learning a continuous-time velocity field guiding a linear interpolation from noise to data:
$\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\,\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}),\quad\text{with the associated velocity field}\quad\mathbf{u}(\mathbf{x}_{t},t)=\boldsymbol{\epsilon}-\mathbf{x}_{0}.$ (2)
A neural network $\hat{\mathbf{u}}_{\theta}(\mathbf{x}_{t},t)$ is trained to approximate this target velocity.
2.2 Architectures for Diffusion Models
Convolutional U-Net with self-attention.
Early latent diffusion models (e.g., Stable Diffusion 1 and 2) rombach2022high primarily employed convolutional U-Net architectures ronneberger2015u , featuring hierarchical downsampling, residual convolutions, and skip connections. Within these architectures, self-attention is occasionally applied to intermediate spatial feature maps to model global dependencies.
@@ 2.2 Architectures for Diffusion Models
Vaswani et al. (vaswani2017attention, ) demonstrated that, rather than performing a single attention operation, it is beneficial to linearly project queries, keys, and values into multiple subspaces and compute attention in parallel. This technique is known as the multi-head attention mechanism.
Specifically, at the $l$-th layer of the denoising network $\hat{\boldsymbol{\epsilon}}_{\theta}$, each attention head $h\in\{1,\dots,H_{l}\}$ applies separate linear projections:
$\mathbf{Q}_{l,h}=\mathbf{Q}_{l}\mathbf{W}_{l,h}^{Q},\quad\mathbf{K}_{l,h}=\mathbf{K}_{l}\mathbf{W}_{l,h}^{K},\quad\mathbf{V}_{l,h}=\mathbf{V}_{l}\mathbf{W}_{l,h}^{V},$ (3)
where $\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l}\in\mathbb{R}^{N\times d}$ are the query, key, and value matrices computed at layer $l$, where $N$ is the sequence length and $d$ is the feature dimension, $\mathbf{W}_{l,h}^{Q},\mathbf{W}_{l,h}^{K},\mathbf{W}_{l,h}^{V}\in\mathbb{R}^{d\times\bar{d}}$ are the head-specific projection matrices, $\bar{d}$ is the dimensionality of each head, and $H_{l}$ is the number of the attention heads in the $l$-th layer. The attention map $\mathbf{A}_{l,h}$ and head output $\mathbf{O}_{l,h}$ are computed as:
$\mathbf{A}_{l,h}=\text{Softmax}\left(\frac{\mathbf{Q}_{l,h}\mathbf{K}_{l,h}^{\top}}{\sqrt{\bar{d}}}\right),\quad\mathbf{O}_{l,h}=\mathbf{A}_{l,h}\mathbf{V}_{l,h}.$ (4)
The outputs from all heads are then aggregated using an output linear projection:
@@
where $\mathbf{W}_{l}^{O}\in\mathbb{R}^{(H_{l}\times\bar{d})\times d}$ is the output projection matrix at layer $l$. For brevity, we omit the layer index $l$ in the remainder of this section when the context is clear:
Multi-head attention allows the model to jointly attend to information from different representation subspaces at different positions. Prior works dosovitskiy2020image ; park2024cross ; gandelsman2023interpreting have sought to interpret and analyze the role of individual attention heads in both language and vision transformers, revealing that heads often capture semantically meaningful and diverse functions.
Diffusion Transformer (DiT).
Transformer-based architectures have played a significant role in recent progress across various machine learning domains devlin2019 ; radford2018 ; dosovitskiy2020 ; chen2020 , offering favorable scalability with respect to model capacity and data size. In the context of diffusion models, the Diffusion Transformer (DiT) peebles2023scalable replaces convolutional structures with pure multi-head attention and demonstrates strong scaling behavior. This has motivated the development of various transformer-based diffusion models chen2023pixart ; chen2024pixart ; chen2024pixartdelta ; nanye2024sit .
Unlike U-Net ronneberger2015u architectures with explicit multi-resolution structures, DiT lacks a coarse-to-fine synthesis hierarchy, making it less clear which layers capture high-level semantic features avrahami2024stable . This structural difference can complicate the application of conventional image editing and perturbation techniques.
More recently, multi-modal attention-based Diffusion Transformers (MM-DiT)esser2024scaling ; blackforestlabs_flux1_dev ; fal_aura_flow have been gaining traction in large-scale text-to-image diffusion. MM-DiT takes as input a concatenated sequence of patchified latent pixels and text embeddings, enabling fully bidirectional interactions between visual and textual modalities. Due to their scalability and performance, MM-DiTs are now used as the backbone in state-of-the-art models such as Stable Diffusion 3 esser2024scaling , FLUX.1 blackforestlabs_flux1_dev .
Unless otherwise noted, all experiments in the main paper use the publicly available Stable Diffusion 3111https://huggingface.co/stabilityai/stable-diffusion-3-medium as the default architecture. Additional results using FLUX.1-Dev [^2] https://huggingface.co/black-forest-labs/FLUX.1-dev are provided in Appendix.
2.3 Diffusion Guidance Mechanisms
Classifier-free guidance (CFG).
Classifier-free guidance ho2022classifier enhances conditional generation without relying on an external classifier. CFG trains the diffusion model to support both conditional and unconditional generation by randomly dropping conditioning tokens during training. At inference, the prediction is interpolated between conditional and unconditional outputs:
@@ 2.3 Diffusion Guidance Mechanisms
where $w$ is the guidance scale, and $\hat{\boldsymbol{\epsilon}}_{\text{cond}},\hat{\boldsymbol{\epsilon}}_{\text{uncond}}$ denote predictions with and without conditioning, respectively.
@@
As attention maps inherently encode spatial and semantic structures cao2023masactrl ; nam2024dreammatcher , perturbing them and guiding the denoising trajectory away from that “bad” prediction can enhance the fidelity or structural quality of the image at test time ahn2024self ; hong2024smoothed ; tero2024 . Analogous to CFG ho2022classifier , attention perturbation guidance extrapolates the original and perturbed outputs:
@@
where $\hat{\boldsymbol{\epsilon}}_{\text{perturbed}}$ denotes the predicted epsilon under attention perturbation. Specifically, this is achieved by modifying the attention maps $\mathbf{A}_{l,h}$ in Eq. 4 during the forward pass, resulting in the perturbed prediction $\hat{\boldsymbol{\epsilon}}_{\text{perturbed}}$.
[…]
Specifically, perturbed-attention guidance (PAG) ahn2024self directly replaces the attention maps in certain layers with an identity matrix $\mathbf{I}\in\mathbb{R}^{N\times N}$:
$\mathbf{A}_{l,h}^{(\text{PAG})}=\mathbf{I},$ (8)
effectively disabling contextual aggregation.
An alternative approach, smoothed energy guidance (SEG) hong2024smoothed indirectly perturbs attention by applying a Gaussian blur $\mathbf{G}_{\sigma}$ to the queries:
$\mathbf{Q}_{l,h}^{(\text{blur})}=\mathbf{G}_{\sigma}\ast\mathbf{Q}_{l,h},\quad\mathbf{A}_{l,h}^{(\text{SEG})}=\text{Softmax}\left(\frac{\mathbf{Q}_{l,h}^{(\text{blur})}\mathbf{K}_{l,h}^{\top}}{\sqrt{\bar{d}}}\right),$ (9)
softening the attention distribution.
The perturbations of PAG and SEG produce aligned but weakened models that still maintain the same architecture. However, existing perturbation guidance methods have typically treated attention maps of a layer as monolithic entities, following the common editing convention Hertz2022PrompttoPromptIE . This overlooks the fact that individual attention heads are often highly specialized and attend to different aspects of the input features vaswani2017attention ; park2024cross ; dosovitskiy2020image ; gandelsman2023interpreting . Moreover, both methods perturb entire layers chosen ad-hoc; no systematic rule exists for where to intervene. We therefore treat a single attention head as the perturbation unit, analyzing its effect in the following section.
@@ 3 Motivation
Figure 1: Motivating example. Each image is generated with PAG ahn2024self , where perturbation is applied to a single attention head within DiTs. Guiding with different perturbed attention heads produces notably distinct results. Results for additional heads are provided in Appendix E.1. All images are generated with the prompt “smiling girl holding a cat, in a flower garden” using stable-diffusion-3-medium. Each row corresponds to a single layer, with different heads perturbed across columns.
Attention perturbation guidance ahn2024self ; hong2024smoothed ; tero2024 steers generation away from weaker predictions by slightly altering the model’s forward pass. Thus, deciding what to weaken becomes critical, but principled methods for determining where in the diffusion network to apply perturbation remain underexplored. As a result, prior works often rely on heuristic selection of perturbation targets, such as the middle block of a U-Net ronneberger2015u , which is known to process high-level semantic information with global self-attention hong2023improving ; ahn2024self .
However, unlike U-Net, transformer-based architectures as in DiTs peebles2023scalable ; chen2024pixart ; esser2024scaling ; blackforestlabs_flux1_dev ; fal_aura_flow lack a coarse-to-fine synthesis structure avrahami2024stable . This absence of an explicit bottleneck makes it challenging to directly identify layers responsible for high-level semantics. Instead, the model distributes semantic processing more uniformly across layers. In this setting, attention, particularly multi-head self-attention vaswani2017attention , plays a central role in modeling global dependencies without convolutions.
Each attention head processes different aspects of the input, and prior works in large language models and vision transformers have shown that heads specialize in capturing distinct semantic attributes vaswani2017attention ; park2024cross ; dosovitskiy2020image ; gandelsman2023interpreting . This suggests that attention heads, rather than entire layers, may serve as more effective units for applying perturbation.
@@ 3.1 Analysis of head-level perturbation guidance
where $\mathbf{A}_{l,h}^{(\text{PAG})}$ denotes the perturbed attention map of head $h$ at layer $l$, and $\mathbf{I}$ is the identity matrix used in PAG ahn2024self . While we illustrate this with PAG, the formulation also applies to other perturbation methods such as SEG hong2024smoothed , with each head perturbed independently. See Appendix B for details.
Figure 2: Generated images from head- and layer-level perturbation guidance. (a) Results of layer-level perturbation guidance, where perturbation is applied to all heads in the layer. (b) Results of head-level perturbation guidance, where each result is obtained by independently applying perturbation to a single head of the layer. Red boxes indicate high-performing heads in terms of PickScore kirstain2023pick . (c) Perturbation guidance using only the high-performing heads identified in (b) yields higher-quality generations across both low-performing layers (L3, L13) and well-performing ones (L8, L12). The prompt “Turkish girl with lantern, dark room” is used.
Intra-layer heterogeneity.
Even when we perturb the attention map of a single layer within diffusion models, it exhibits highly polysemantic attentive behavior, as the computation in individual heads occurs independently and diversely elhage2021mathematical ; park2024cross . Therefore, applying the same perturbations across all heads in a layer fails to account for this intra-layer diversity, potentially disrupting the output from the guidance and leading to unintended side effects. As shown in Fig. 2 (a), some heads from supposedly poor layers still produce high-quality results. This suggests that overall quality may degrade due to the effects of a few heads dominating the outcome. To test this, we filtered and perturbed only high-scoring heads based on preference scores. The resulting samples (Fig. 2 (c)) show clear improvements over standard layer-level guidance. Moreover, even well-performing layers selected through manual layer-level search can include suboptimal heads, and removing them further improves quality (L8). These results demonstrate the inefficiency of coarse layer-level guidance. Based on these results, we suggest analyzing head-level guidance and its properties.
@@
Perturbing more heads strengthens the guidance effect but can also introduce undesirable artifacts such as oversaturation or oversimplification. As shown in Fig. 4 (a), adding more perturbed heads results in over-perturbation, leading to over-smoothed or overly simplified output. Fig. 4 (b) further illustrates this effect quantitatively: the quality metric (FID kirstain2023pick ) using 1K prompts from the MS COCO dataset begins to degenerate as more heads are added after some threshold. This suggests that, in addition to the choice of perturbation method, the number of perturbed heads itself is a key factor for controlling the overall perturbation strength.

### v4

# Donghoon Ahn∙1   Jiwon Kang∙1 Sanghyun Lee∘1  Minjae Kim∘2   Jaewon Min1   Wooseok Jang1

@@ 1 Introduction
Diffusion models [19, 54, 53, 56, 58, 11, 45, 42] and flow-matching models [34, 37, 14] have gained great popularity in visual generation tasks, including images [45, 42, 41, 5, 14], videos [21, 18, 3, 63], 3D [43, 32, 60, 50], and 4D content [51, 64, 61]. The key behind the success of diffusion models lies in classifier-free guidance (CFG) [20], which substantially enhances image generation quality during conditional inference. Despite its effectiveness, CFG has two key limitations. First, it applies only to conditional generation, limiting its applicability in unconditional settings such as inverse problems [25, 8, 55, 9]. Second, CFG often reduces sample diversity and leads to over-saturated or overly simplified outputs [47, 28, 48, 26].
To address the limitations of CFG in unconditional scenarios, alternative guidance strategies [1, 23, 22, 26, 24] have been proposed. These methods steer the denoising trajectory by perturbing the input or the model itself, or by training a weaker model, thereby guiding samples away from low-quality regions and toward the high-quality data manifold. Among these strategies, attention-layer perturbation approaches [1, 22] continue to be a practical and widely explored strategy, as they can generate well-aligned weak models without additional training. Some papers explain these guidance methods from an energy perspective [22] and draw connections to Hopfield networks [10, 44], which empirically work well with specific blocks (medium blocks of U-Net [46] architecture) of attention.
However, there is still limited understanding of where perturbations should be applied. In particular, Diffusion Transformer (DiT) [41] architectures lack localized blocks responsible for global semantics, unlike U-Nets [46], and instead distribute this functionality more evenly across all layers [2]. This structural difference makes it even more critical to carefully select perturbation targets to achieve effective guidance.
To determine suitable perturbation targets, we first consider the underlying structure of DiT. In DiT, multi-head self-attention [59] plays a central role in modeling global dependencies. Each head attends to different aspects of the input, and prior work has shown that heads often specialize in distinct semantic features [59, 40, 12, 16]. This suggests that attention heads, rather than entire layers, may serve as more precise and effective targets for perturbation.
@@
Building on these observations, we (i) analyze key properties of head-level perturbation guidance, including composability and controllability, (ii) propose HeadHunter, a systematic framework for retrieving heads that align with arbitrary objectives, and (iii) introduce SoftPAG, a variant of PAG [1] that enables fine-grained, continuous control over guidance strength, mitigating over-smoothness and oversimplification caused by overly aggressive perturbations.
@@ 2 Preliminaries
Diffusion and flow matching models.
Diffusion models [19, 52, 57] define a generative process as iterative denoising of a sample drawn from a Gaussian prior. The forward process gradually perturbs a clean data point $\mathbf{x}_{0}\sim p_{\text{data}}$ through the noise schedule $(\alpha_{t},\sigma_{t})$, as $\mathbf{x}_{t}=\alpha_{t}\mathbf{x}_{0}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $t\in[0,1]$. A neural network $\hat{\boldsymbol{\epsilon}}_{\theta}(\mathbf{x}_{t},t)$ learns to predict the added noise $\boldsymbol{\epsilon}$, enabling reverse-time sampling from noise to data. On the other hand, flow matching [35, 36] provides a deterministic alternative by learning a continuous-time velocity field that guides a linear interpolation from noise to data: $\mathbf{x}_{t}=(1-t)\mathbf{x}_{0}+t\,\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, and the associated velocity field is $\mathbf{u}(\mathbf{x}_{t},t)=\boldsymbol{\epsilon}-\mathbf{x}_{0}$. A neural network $\hat{\mathbf{u}}_{\theta}(\mathbf{x}_{t},t)$ is trained to approximate this target velocity.
@@
Vaswani et al. [59] introduced the multi-head attention mechanism, which projects queries, keys, and values into multiple subspaces and computes attention in parallel. At the $l$-th layer of the denoising network $\hat{\boldsymbol{\epsilon}}_{\theta}$, each head $h\in\{1,\dots,H_{l}\}$ applies distinct linear projections:
$\mathbf{Q}_{l,h}=\mathbf{Q}_{l}\mathbf{W}_{l,h}^{Q},\;\mathbf{K}_{l,h}=\mathbf{K}_{l}\mathbf{W}_{l,h}^{K},\;\mathbf{V}_{l,h}=\mathbf{V}_{l}\mathbf{W}_{l,h}^{V},$ (1)
where $\mathbf{Q}_{l},\mathbf{K}_{l},\mathbf{V}_{l}\in\mathbb{R}^{N\times d}$ are the query, key, and value matrices, $\mathbf{W}_{l,h}^{Q,K,V}\in\mathbb{R}^{d\times\bar{d}}$ are the head-specific projections, $\bar{d}$ is the per-head dimensionality, and $H_{l}$ is the number of attention heads in layer $l$. The attention map and output for each head are
$\mathbf{A}_{l,h}=\text{Softmax}\!\left(\frac{\mathbf{Q}_{l,h}\mathbf{K}_{l,h}^{\top}}{\sqrt{\bar{d}}}\right),\quad\mathbf{O}_{l,h}=\mathbf{A}_{l,h}\mathbf{V}_{l,h},$ (2)
and all head outputs are concatenated and projected:
@@
where $\mathbf{W}_{l}^{O}\in\mathbb{R}^{(H_{l}\bar{d})\times d}$ is the output projection matrix. Multi-head attention enables the model to capture diverse semantic relations across tokens [12, 40, 16].
Classifier-free guidance.
Classifier-free guidance (CFG) [20] enhances conditional generation by extrapolating predictions from conditional and unconditional models:
@@
where $w$ is the guidance scale, and $\hat{\boldsymbol{\epsilon}}_{\text{cond}}$, $\hat{\boldsymbol{\epsilon}}_{\text{uncond}}$ denote the predicted noise with and without conditioning, respectively.
@@
As attention maps encode spatial and semantic structures [4, 39], perturbing them can guide the denoising trajectory away from suboptimal predictions, improving structural fidelity at test time [1, 22, 26]. Analogous to classifier-free guidance [20], attention perturbation guidance extrapolates between original and perturbed predictions:
@@
where $\hat{\boldsymbol{\epsilon}}_{\text{perturbed}}$ is obtained by modifying attention maps $\mathbf{A}_{l,h}$ in Eq. 2 during the forward pass.
Let $\mathcal{L}$ denote the perturbed layers and $H_{l}$ the number of heads in layer $l$. Perturbations are applied to $\mathbf{A}_{l,h}$ for all heads $h\in\{1,\dots,H_{l}\}$ and $l\in\mathcal{L}$. In perturbed-attention guidance (PAG) [1], the attention maps are replaced with the identity matrix, disabling contextual aggregation:
$\mathbf{A}_{l,h}^{(\text{PAG})}=\mathbf{I}\in\mathbb{R}^{N\times N}.$
@@ 3 Motivation
Figure 1: Motivating example. Each image is generated with PAG [1], where perturbation is applied to a single attention head within DiTs. Guiding with different perturbed attention heads produces notably distinct results. Results for additional heads are provided in Appendix E.1. All images are generated with the prompt “smiling girl holding a cat, in a flower garden” using stable-diffusion-3-medium. Each row corresponds to a single layer, with different heads perturbed across columns.
Attention perturbation guidance [1, 22, 26] steers generation away from weaker predictions by slightly altering the model’s forward pass. Thus, deciding what to weaken becomes critical, but principled methods for determining where in the diffusion network to apply perturbation remain underexplored. As a result, prior works often rely on heuristic selection of perturbation targets, such as the middle block of a U-Net [46], which is known to process high-level semantic information with global self-attention [23, 1].
However, unlike U-Net, transformer-based architectures as in DiTs [41, 5, 14, 29, 15] lack a coarse-to-fine synthesis structure [2]. This absence of an explicit bottleneck makes it challenging to directly identify layers responsible for high-level semantics. Instead, the model distributes semantic processing more uniformly across layers. In this setting, attention, particularly multi-head self-attention [59], plays a central role in modeling global dependencies without convolutions.
Each attention head processes different aspects of the input, and prior works in large language models and vision transformers have shown that heads specialize in capturing distinct semantic attributes [59, 40, 12, 16]. This suggests that attention heads, rather than entire layers, may serve as more effective units for applying perturbation.
@@
Figure 2: Generated images from head- and layer-level perturbation guidance. (a) Results of layer-level perturbation guidance, where perturbation is applied to all heads in the layer. (b) Results of head-level perturbation guidance, where each result is obtained by independently applying perturbation to a single head of the layer. Red boxes indicate high-performing heads in terms of PickScore [27]. (c) Perturbation guidance using only the high-performing heads identified in (b) yields higher-quality generations across both low-performing layers (L13) and well-performing ones (L12). The prompt “Turkish girl with lantern, dark room” is used.
@@ 3.1 Analysis of head-level perturbation guidance
where $\mathbf{A}_{l,h}^{(\text{PAG})}$ denotes the perturbed attention map of head $h$ at layer $l$, and $\mathbf{I}$ is the identity matrix used in PAG [1]. While we illustrate this with PAG, the formulation also applies to other perturbation methods such as SEG [22], with each head perturbed independently. See Appendix F for details.
Inefficiency of layer-level guidance due to intra-layer diversity.
Even when we perturb the attention map of a single layer within diffusion models, it exhibits highly polysemantic attentive behavior, as the computation in individual heads occurs independently and diversely [13, 40]. Therefore, applying the same perturbations across all heads in a layer fails to account for this intra-layer diversity, potentially disrupting the output from the guidance and leading to unintended side effects. As shown in Fig. 2, some heads from supposedly poor layers (L13 of Fig. 2 (a)) still produce high-quality results (Fig. 2 (b)). This suggests that overall quality may degrade due to the effects of a few heads dominating the outcome, indicating further room for improvement. To test this, we filtered and perturbed only high-scoring heads based on preference scores. The resulting samples (Fig. 2 (c)) show clear improvements over standard layer-level guidance. These results demonstrate the inefficiency of coarse layer-level guidance and motivate us to analyze head-level guidance and its properties.
@@
Perturbing more heads strengthens the guidance effect but can also introduce undesirable artifacts such as oversaturation or oversimplification. As shown in Fig. 4 (a), adding more perturbed heads results in over-perturbation, leading to over-smoothed or overly simplified output. Fig. 4 (b) further illustrates this effect quantitatively: the quality metric (FID [27]) using 1K prompts from the MS COCO dataset begins to degenerate as more heads are added after some threshold. This suggests that, in addition to the choice of perturbation method, the number of perturbed heads itself is a key factor for controlling the overall perturbation strength.